OpenStudy (anonymous):

please help me??what is the solution 2x^2+x+2=0

OpenStudy (kc_kennylau):

This is just a simple quadratic :)

OpenStudy (kc_kennylau):

(and the question is wrong: you can't have real answers for that question)

OpenStudy (anonymous):

one of the answers is x=-1+i squareroot15/2

OpenStudy (kc_kennylau):

yep according to wolfram alpha you'd better check if your teacher has put this down wrongly

OpenStudy (anonymous):

alright can you help me with a simple one?

OpenStudy (kc_kennylau):

like?

OpenStudy (anonymous):

the length of the sides of a rectangular window have a ratio 1.6 to 1. the area of the window is 2,560 square inches. what are the dimensions of the window?

OpenStudy (kc_kennylau):

let the short side be x. the long side is therefore 1.6x.

OpenStudy (kc_kennylau):

1.6x^2=2560

OpenStudy (kc_kennylau):

i suppose you know how to do then? :)

OpenStudy (anonymous):

no i dont

OpenStudy (kc_kennylau):

1.6x^2=2560 16x^2=25600 x^2=1600 x=40

OpenStudy (anonymous):

40 inches by ??

OpenStudy (kc_kennylau):

by 40*1.6=64 inches sorry for not replying

OpenStudy (kc_kennylau):

you can use @kc_kennylau next time

OpenStudy (anonymous):

thank you:) alright

OpenStudy (kc_kennylau):

no problem :)

OpenStudy (anonymous):

A football is kicked into the air from an initial height of 4 feet. The heigh in feet of the football avoce the ground is given by s(t)=-16t^2+50t+4 where t in time is seconds and t>=O. At what time will the football be 25 feet above the ground? A)3.625 seconds B)3.20 seconds C)0.5 seconds or 3.625 seconds D)0.5 seconds or 2.625 seconds

OpenStudy (anonymous):

@kc_kennylau

OpenStudy (kc_kennylau):

quadratics again :)

OpenStudy (anonymous):

lol

OpenStudy (kc_kennylau):

-16t^2+50t+4=25

OpenStudy (anonymous):

im still confused

OpenStudy (kc_kennylau):

this is just quadratic :)

OpenStudy (anonymous):

never mind i got it

OpenStudy (kc_kennylau):

good :)

OpenStudy (anonymous):

Suppose a parabola has an axis of symmetry at x = –7, a maximum height of 4, and also passes through the point (–6, 0). Write the equation of the parabola in vertex form.

OpenStudy (kc_kennylau):

ok i don't know much about parabola sorry

OpenStudy (anonymous):

its cool thanks

OpenStudy (kc_kennylau):

maybe you can ask others? :)